Title of Investigation: MRI: Development of Thermal Infrared Spectrograph for the 1.6 Meter Solar Telescope in Big Bear

Principal Investigator: Roy Coulter (BBSO/NJIT)

Co-Principal Investigators: Matt Penn (NSO/T), Tom Ayres (CU) and Philip Goode (BBSO/NJIT)

Proposal Summary

Early in 2009 at Big Bear Solar Observatory (BBSO), first light observations were made with our 1.6 m clear aperture, off-axis telescope - the NST (New Solar Telescope). The NST will be the most capable, largest aperture solar telescope in the US until the 4 m off-axis ATST (Advance Technology Solar Telescope) comes on-line in the middle of the next decade. Here we propose, in collaboration with the Matt Penn of the National Solar Observatory (NSO) and Thomas Ayres of the University of Colorado (CU), to build the Cooled InfraRed Spectrograph (CIRS) spanning 1.0 to 5.0 μm for the NST. This system is essential for the NST to achieve its scientific potential of probing for the first time the fundamental scale of the Sun’s dynamic magnetic field, the origin of space weather, while providing valuable experience and risk reduction for the ATST project.

???SUMMARY TEXT FOR SCIENCE AND SPECTROGRAPH???

The principal scientific objectives for CIRS equipped NST are ???FINISH THIS???

In sum, we directly address the two criteria required in an NSF proposal:

(1) Intellectual merit: CIRS will enable the NST to provide the community unique (???HOW UNIQUE COMPARE TO WHAT CAN BE DONE AT THE MCMATH???) data on the solar magnetic field illuminating the fundamental character of our dynamic star, while probing the origins of space weather. Over the next decade, NSO will build and begin operating the ATST, the flagship telescope for the solar community. CIRS will be a significant testbed for the more complicated IR spectrograph planned for the ATST. After the ATST is on-line, the NST will continue to be an essential telescope in making campaign-style observations.

(2) Broader Impact: The development of the proposed CIRS system continues our decade long tradition of training the next generation of scientists who build instruments. The resulting data will be used in many PhD theses. Our data will be open to the scientific community, and a substantial portion of observing time on the NST will be open to the community.

A. Project Description

A.1 Overview

Spectral observations of the photosphere and chromosphere utilizing the largely unexplored near-infrared (NIR) are crucial for solving many outstanding problems in solar astronomy. In particular, high spatial resolution polarimetry at NIR wavelengths is essential and requires large aperture telescopes. Here, we propose for funding to implement a cyrogenically Cooled IR Spectrograph (CIRS) operating from 1 to 5 μm for the New Solar Telescope (NST) for observations at the Nasmyth bench. Firstlight observations were made in January 2008 at Big Bear Solar Observatory (BBSO) with the new 1.6 m clear aperture, off-axis telescope - the NST. The NST will be the largest aperture solar telescope in the US, and perhaps the world, until the 4 m off-axis ATST (Advanced Technology Solar Telescope) comes on-line in the middle of the next decade. In addition to its tremendous scientific potential, the NST will be the testbed for many essential components and technologies of the ATST including CIRS. Primarily we propose here, in collaboration with Matt Penn of the National Solar Observatory (NSO) and Thomas Ayres of the University of Colorado (CU), to build, which is essential for the NST to reach its scientific potential.

A.2 Schematic of the NST

BBSO has removed its old 0.6 m aperture telescope (donated to the Los Angeles Amateur Astronomy Society) and the alignment of the modern, off-axis 1.6 m clear aperture instrument (Goode et al., 2003) has been finished and first light has been attained at the Nasmyth focus, which will be the location of the thermal IR spectrograph proposed here. The new telescope offers a significant improvement in ground-based high angular resolution capabilities, and will substantially enhance our efforts to understand (???WHAT EVER WILL BE PRO-POSED IN NEXT SECTION???), as sketched in Section A.5. The NST project is described in detail in \underline{http://www.bbso.njit.edu/newtelescope} .

The NST is configured as an off-axis Gregorian system consisting of a parabolic primary, prime fo-

Figure 2: The NST equatorial mount, mirror cell, counterweights (for right ascension balance), and tower to secondary.???UPDATE???

cus field stop and heat reflector (heat-stop), elliptical secondary and diagonal flats, see Figure 1. The primary mirror (PM) is 1.7 m with a clear aperture of 1.6 m with a final figure residual error of 16 nm rms. The PM was figured by Steward Observatory Mirror Lab. The focal ratio of the PM is f/2.4, and the final ratio is f/50. The 180′′circular opening in the field stop defines a 120′′×120′′maximal square field of view (FOV) of the NST. The working wavelength range will cover 0.4 to 1.7 μm in the Coudé Laboratory one floor beneath the telescope, and all wavelengths including far infrared at the Nasmyth focus on an optical bench attached to the side of the telescope structure. Here we are proposing for a spectrograph at the Nasmyth. An off-axis design was chosen principally because of its vastly reduced stray light, since there is no central obscuration, which reduces the telescope’s MTF (Modulation Transfer Function) at high spatial frequency.

We have designed an open telescope structure with an equatorial mount, as illustrated in Figure 1. A key design issue for a large aperture, solar telescope is the thermal control system to keep the temperature of the mirrors very close to, or even slightly below that of the ambient air. The dome employs a windgate and exhaust system to control the airflow from the wind helping to maintain the same temperature inside and outside the dome, and remove concentrations of heat in and around the optical paths. In fact, we have measured the temperatures along the optical path from the telescope to the Coudé Laboratory (which is thermally controlled to 0.5 C). When the vent gates and exhaust system are turned on, the temperature all along the optical path becomes precisely uniform (A. Coulter, 2008, private communication). In addition, we have implemented a closed cycle, chilled air knife system as part of the OSS (Optical Support Structure) to limit so-called “mirror seeing” by sweeping away turbulent cells and directly cooling the PM. What matters for CIRS is the feed to the Nasmyth bench, which also has an isothermal light path. The OSS and active support mirror cell were built and tested by DFM, Engineering in Longmont, CO. The PM is supported by 36 actuators that can bend out low-order aberrations, such as those due to gravity and/or thermal effects. Its shape will be controlled by a wavefront

Figure 1: Schematic of the 1.6 m off-axis, open NST (top) and its optical layout (bottom). On the top part, the heat stop (HS) resides at the Gregorian focus with the SM mounted on a PI hexapod. The wavefront sensing system (WFS) resides immediately before M3 (along with the polarization optics for magnetic field measurements). M4 sends the light to the Coudé Room below. The HS, SM, M3 and M4 are not apparent in the figure. If M4 is flipped the light goes to the Nasmyth Bench where the spectrograph will reside (???REPLACE UPPER DRAWING WITH ONE HAVE A SCHEMATIC SPEC-TROGRAPH ON THE NASMYTH???) for which the light path is shown in red. This is where the first light observations occurred. The design was driven by competing needs for rigidity near the SM and an open light path. For more details, see \underline{http://www.bbso.njit.edu/newtelescope} .

sensing system (\underline{http://www.wavefrontsciences.com} ), which resides immediately before M3. First light observations of the Sun have been made at the Nasmyth focus, and image is shown in Figure 3.

Figure 3: A diffraction-limited image obtained on the NST at the Nasmyth focus (???NEED REAL FIGURE WITH R0 TO DEMONSTRATE KOLMOGOROV SEE-ING, ETC???).

The new, larger dome, required to house the NST, is fully operational and the installed telescope is shown in Figure 2. The elliptical secondary mirror (SM) was silvered, and was been tested by both SORL (Space Optics Research, which figured the SM) and Steward before its installation on the NST PI hexapod. The latest version of the BBSO telescope headquarters control system software has been written in BBSO and tested at DFM. While this software is not essential for early observations, it would make them somewhat less cumbersome.

A.3 Strehl Ratio of the Beam Feeding the Spectrograph

Ideally, we would like the Nasmyth Bench spectrograph to be fed with diffraction limited light over the entire 1.0-5.0 μm range of operation. This spectral range does not require full adaptive optics correction because the terrestrial atmosphere is less turbulent in the IR than in the visible, so tip-tilt corrections alone are sufficient for most observations. Given the see-

3 ing at BBSO and the limitations of tip-tilt corrections, the diffraction limit will almost never be achieved at 1.0 μm, while for wavelengths beyond 2.5 μm, diffraction limited observations will be the norm.

The Strehl, generally used to quantify the performance of an adaptive optics (AO) system, is the ratio of the maximum intensity in the corrected image in the detector plane to that from a theoretical, perfect imaging system operating at the diffraction limit. Experience, at the National Solar Observatory at Sacramento Peak (NSO/SP) and elsewhere, from existing solar AO systems as well as detailed performance modelling of corrected observations using MHD simulations of solar magnetic fields show that a Strehl of S ≥ 0.3 is required to perform the desired quantitative analysis of the small-scale magnetic elements. So, under what circumstances will this Strehl be achieved for light fed to our planned spectrograph by a tip-tilt system?

A simple way of quantifying the image quality after tip-tilt correction is by computing the residual wavefront error variance, σ^{2} = 0.134(D/r_{0})^{\frac{5}{3}}. From, this we can calculate the Strehl of a long exposure image in the usual Marechal approximation (for more detailed discussion see Christou et al, 2006a and Christou et al., 2006b, and references therein), where S = e^{−σ^{2}} . Further, D and r_{0} are the telescope aperture diameter and r_{0} is the Fried parameter. Often, the Fried parameter, r_{0}, is used to quantify the seeing conditions at a site. r_{0} is the coherence length of the atmospheric turbulence. Without image correction, for telescope apertures larger than r_{0}, resolution is seeing limited. The Fried parameter is wavelength dependent (r_{0} ∝ λ^{\frac{6}{5}} for the generally-assumed Kolmogorov turbulence). The BBSO site has been well-studied within the ATST site survey (see \underline{http://atst.nso.edu/site}for details). Over the two years of seeing measurements, BBSO has been shown to be characterized by steady seeing all day long with an average of Fried parameter of about 6.5 cm for light of λ=0.5μm. About onequarter of the time, r_{0} ≥10 cm all day long. The origin for the steady, good seeing is the well-documented lack of ground-level atmospheric turbulence, which is the result of the NST dome being located in the middle of Big Bear Lake. The lake is almost always cooler than the air and, thus, provides a natural inversion that suppresses the turbulence during our typical observing conditions in which a constant wind comes from the west over two miles of open lake water.

Using the Marechal approximation and assuming r_{0} ∼10 cm at λ=0.5μm in all cases in this paragraph, we calculate that at 1 μm r_{0} ∼25 cm, and we have S=0.03, which means we are seeing limited with a seeing disk of about 0′′.5. At 1.6 μm, where the Sun has its opacity minimum and there are strong magnetic lines, we have r_{0} ∼40 cm and a tip-tilt corrected S=0.26, which is close to the 0.3 that experience has taught us is sufficient for diffraction limited imaging that would be 0′′.20 at 1.6 μm for the spectrograph. At 2.5 μm, r_{0} ∼70 cm and S ∼ 0.7. For all wavelengths red of 2.5 μm, even the everyday seeing, after tip-tilt correction, is sufficient to feed diffraction limited light to the spectrograph. In general, we plan to use our real-time speckle reconstruction system as a daily guide on the wavelengths feeding the spectrograph. Some particular wavelengths of interest in the 1-5 μm band are listed in Table 1, along with their significant properties and the science tasks that can be attacked with the spectrograph. Another advantage at longer-than-visible wavelengths is the appreciably greater Zeeman splitting, which enables us to separate the true magnetic field strength and filling factor from the magnetic flux that is measured in the visible.

A.4 Zeeman Splitting in the NIR

Observations in the NIR have advantages associated with the relatively larger Zeeman splitting (Zeeman splitting increases quadratically with wavelength) in that regime and are vital for measurements of weak magnetic field strengths. In that way, visible and NIR observations complement each other.

The Zeeman splitting, δλ, is given by δλ ∝ gBλ^{2}, or more typically the fractional Zeeman splitting δλ/λ ∝ gBλ = λ(gB), where λ is the wavelength, g is the Landé g-factor and B is the magnetic field strength. Further, (gB) is called the “Magnetic Sensitivity” for purposes of comparing the utility of various IR lines.

A.5 Scientific Motivation: Photospheric and Chromospheric Magnetic Fields

Many fundamental properties of our dynamic star will finally come within our grasp with the tip-tilt corrected IR light feeding the NST Thermal IR Spectrograph, CIRS.

Table 1: IR Bands of Interest for the Nasmyth Spectrograph.